QUESTION 3 (20 marks) Draw the shear and moment diagrams for the following beam. 2 kN/m 2 kN/m 18 kNm 3 m3 m 6 kNhi, this is settling an argument so please get it right:)
Answer #1

solution:

Solution :- 2KN/m 2KNIM 6KN 3m — * 3m - GKN SED 12 KNMI 8 MD 2 (3-2) 2KN/m i In BC (a from c) Sn = (2+ 2/3 (3-2)] BK3- 12 AtMa = ( 213(3-1) + 2] x [4+ 213 (3-2)] 2+2/3 (3-2) At a=0 Mc = - 1 (2+2)x0= 0 kNm At x=3m, MB=-1/2 x 2 x 33 x 3x (4*0 ato = -6The first denivative of the sf equation at any point gives the loading detail at that point → The load is hon zero at A and c